OpenStudy (anonymous):

For the function f(x)=2-3x, find f(4) A. -10 B. -5 C. 12 D. 14 Help please? :)

OpenStudy (anonymous):

f(4)=2-3*4=2-12=-10

OpenStudy (imtiaz7):

A is your answer

OpenStudy (hitaro9):

Alright, so functions are essentially tools where you put something in, the function has a set of rules that modifies what you put in, and returns something new. So in this, the set of rules says (We subtract, from 2, 3 times what you put in) So you put in 4 (insert it in the x slot) f(4) = 2 - 3(4) f(4) = 2 - 12 f(4) = -10 -10 is what you get out, and is your answer. So pick A.